OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 4 sin2x - 4 sin x + 1 = 0 I'd like to know the identities used and how you came to know which one to use.

OpenStudy (anonymous):

(4 sin^2x) - 4 sin x + 1 = 0

OpenStudy (amistre64):

youd use the quadratic equation ....

OpenStudy (amistre64):

let u = sin(x)

OpenStudy (amistre64):

4 sin^2 x - 4 sin x + 1 = 0 4 u^2 - 4u + 1 = 0 solve for u to determine sin(x)

OpenStudy (anonymous):

Wow... I feel dumb. Thank you very much!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

Where do I go from (2u-1)^2 = 0?